OpenStudy (anonymous):

One side of a square is 5n inches long. Another side is n + 16 inches long. Write an equation by setting these equal and solving for n. How long is a side of the square? One point for finding n and one point for finding the length of the side.

OpenStudy (anonymous):

All sides of a square are equal, therefore, 5 n =n+16 4n = 16 subtract n from both sides n=4 divide both sides by 4, value of n 5*4 = 4+16 = 20 length of side